If NP2=90 , then find n

Answer by Edwin McCravy(20054)


NP2 = 90
NP2 means N(N-1)

N(N-1) = 90

N(N-1) = 90

N(N-1) = 90

N2 - N = 90

N2 - N - 90 = 0

(N-10)(N+9) = 0

N-10=0;  N+9=0
   N=10;   N=-9

Ingnore negative answer.

Check:

10P2 = 10(9) = 90

Edwin